form 9x+3y-6 = 0 Of Ax+By=C
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52781)

.

    A = 9

    B = 3 

    C = 6

Solved, answered and explained. And completed.



Answer by greenestamps(13200)



Take the time to be polite and actually ask a question, preferably in full sentences....

I THINK what you want to do is put the given equation in the specified form, Ax+By=C.

If that is it, then all you need to do is get the constant on the right side by adding 6 to both sides of the given equation.
   9x+3y-6 = 0
        +6  +6
  ------------
   9x+3y   = 6

ANSWER: 9x+3y = 6